Fermi National Laboratory


Accelerator Update for January 25, 2006 - January 27, 2006

Wednesday January 25

The day shift began with Tevatron (Tev) off due to TEL problems, the Recycler Ring (RR) stash at 4.4 E10, Antiproton Source (Pbar) down because of F23 breaker, and with MiniBooNE and NuMI taking beam.

At 9:18 AM, F23 breaker was reset.

Pbar returns to stacking at 9:58 AM.

TEL work was completed at 10:21 AM.

Beam to MIPP at 10:54 AM.

Tev is ramping at 12:02 PM, they will begin studies.

Operations put the Recycler into shot setup at 2:24 PM to stash antiprotons.

Beam returns to MIPP at 2:50 PM, Pbar remains off to investigate a potential problem with the lithium lens.

Access into Pbar rings and transport began at 4:31 PM, NuMI and MiniBooNE are off due to the access,

Pbar transport access is complete at 4:59 PM, beam returns to NuMI and MiniBooNE.

Meson Test (MT) begins taking beam at 6:00 PM.

At 8:11 PM, it appears that the lithium lens for PBar is shorted to ground, experts will be back at 6:30 AM to continue working on the lens. No Pbar beam until this is repaired.

Thursday January 26

The midnight shift began with Tevatron in studies mode, the Recycler Ring (RR) stash at 29.5 E10, Antiproton source off, and with MiniBooNE, NuMI, MIPP and MT taking beam.

At 3:34 AM, MT turned off for the night.

At 6:00 AM, Transport, NuMI and MiniBooNE were turned off to prepare for accesses.

Transport access begins at 7:20 AM.

NuMI, MiniBooNE and Meson are also going into access.

Tev studies were completed at 10:11 AM.

D0 goes into access at 11:24 AM, to investigate a grounding problem.

Transport access was completed at 9:17 PM.

Beam to NuMI at 9:31 PM.

D0 finished their access at 10:22 PM.

Beam to MiniBooNE at 11:00 PM

Pbar returns to stacking at 11:01 PM.

Tev resumes its studies at 11:42 PM.

Friday January 27

The midnight shift began with Tevatron in studies mode, the Recycler Ring (RR) stash at 22.1 E10, Pbar stacking, and with MiniBooNE and NuMI taking beam.

MIPP resumed taking beam at 12:19 AM.

The Tev quenched (B4) at 1:39 AM, investigating the cause of the quench.

At 6:28 AM, power supply MI50U tripped off due to a water leak at the supply, experts are investigating. This keeps beam off to stacking, NuMI and Meson.

At 7:48 AM, the B0 cooldown has pretty much stalled, the Tev is being prepared for access to add some turbo pumps to the vacuum.

The Plan for Friday and the Weekend

Get MI50U on and resume stacking, resume beam to NuMI and MIPP. Get the Tev on and resume studies until we have sufficient Pbars for a Tev store. Return to normal operations of stacking, stashing and storing.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 01/30/2006    email Fermilab